OpenStudy (anonymous):

Find the slope of lines containing (1,6) and (10,-9)

OpenStudy (anonymous):

-5/3

OpenStudy (anonymous):

m = (y2 - y1) / (x2 - x1) <------------SLOPE FORMULA m = (-9 - 6) / (10 - 1) m = -15/9 m= -5/3

OpenStudy (anonymous):

-5/3 is correct!